# Phase Retrieval by Quaternionic Reweighted Amplitude Flow on Image Reconstruction

### v1

@@ Abstract
[Ren1]organization=School of Computing and Information Technology, Great Bay University,addressline=Songshanhu international community, Songshanhu District, city=Dongguan, postcode=523000, state=Guangdong, country=China
@@
[Ren2]organization=Tsinghua Shenzhen International Graduate School, Tsinghua University,addressline=Xili University Town, Nanshan District, city=Shenzhen, postcode=518071, state=Guangdong, country=China
[…]
[Pan]organization=School of Mathematical Sciences, Tianjin Normal University,addressline=Binshui No.393, Xiqing District, city=Tianjin, postcode=300387, state=Tianjin, country=China
@@ 1 Introduction
Over the past three decades, the use of hypercomplex numbers, such as quaternions, split quaternions, octonions, and Clifford algebras, has garnered significant attention in the representation of multidimensional signals, particularly in addressing challenges in signal and image processing, see for instance [1, 3, 6, 7]. This methodology has found wide applications in several domains, including optical imaging, array processing, wireless communications, filtering, and neural networks, see e.g., [10, 11, 12, 13, 22]. A key advantage of this approach lies in their ability of hypercomplex numbers to exploit intrinsic correlations across different signal dimensions, often resulting in compact, elegant, and interpretable frameworks for managing the geometric properties of multidimensional signals and images. This benefit arises from the algebraic structure of hypercomplex systems, which allow for both addition and multiplication, in contrast to the usual vector spaces that permit only addition and scalar multiplication. Very recently, hypercomplex techniques were applied to tackle the problem of phase retrieval in high-dimensional settings, leveraging these algebraic advantage. We refer to the recent survey [2] for the exciting new developments.
1.1 Phase retrieval and algorithm development
The original phase retrieval problem entails the reconstruction of a signal from the magnitude of its Fourier transform or the magnitude of its short-time Fourier transform, see e.g., [14, 15]. This problem has been extended to encompass general measurements. Formally,
Problem 1.1.
The generalized phase retrieval problem seeks to recover a real or complex-valued signal $\boldsymbol{x}\in\mathbb{R}^{d}$ or $\mathbb{C}^{d}$ from its amplitude-only measurements given by
$b_{j}=\left|\langle\boldsymbol{\alpha}_{j},\boldsymbol{x}\rangle\right|,\qquad j=1,\ldots,n,$
where $\boldsymbol{\alpha}_{j}$ are the measurement vectors taken from a known measurement matrix $\boldsymbol{A}\in\mathbb{R}^{n\times d}$ or $\mathbb{C}^{n\times d}$.
This problem arises across various scientific and engineering areas, including diffractive imaging, X-ray crystallography, astronomy, and radar waveform design [14]. Significant efforts have been dedicated to developing algorithms for conventional real and complex phase retrieval problems, which can be broadly classified into convex and non-convex approaches. While convex optimization methods (see e.g., [18, 19]) usually offer strong theoretical recovery guarantees and demonstrate strong performance, they often encounter high computational complexity in large-scale problems and may require large oversampling ratios for exact recovery. On the other hand, efficient non-convex approaches have emerged. Notably, Netrapalli et al. introduced the AltMinPhase algorithm in 2013 in [26]. More recently, Candès et al presented the Wirtinger Flow (WF) algorithm in their seminal work [20], which guarantees signal recovery through a simple gradient descent approach. Specifically, the WF algorithm resolves the Phase Retrieval problem via two key steps:
  1. 1.
    Obtaining a guess initialization using spectral method;
  2. 2.
    Iteratively refining this initial estimate through a gradient descent scheme to solve the intensity-based model
    $\min_{\boldsymbol{x}\in\mathbb{F}^{\,d}}F(\boldsymbol{x})=\frac{1}{n}\sum_{j=1}^{n}\left(\left|\langle\boldsymbol{\alpha}_{j},\boldsymbol{x}\rangle\right|^{2}-b_{j}^{2}\right)^{2},$ (1.1)
where $\mathbb{F}\in\{\mathbb{R},\mathbb{C}\}$. The success of gradient descent in WF is guaranteed by the favorable benign geometric landscape established in [31] by Sun et al. The method (WF) is further enhanced by the truncated Wirtinger flow (TWF) in [21], which employs a Poisson loss function, significantly reducing both the sample complexity to $\mathcal{O}(d)$ and the computational complexity to $\mathcal{O}(nd\log 1/\varepsilon)$.
Moreover in the real field $\mathbb{R}$, WF is further refined by addressing the following amplitude-based model
$\min_{\boldsymbol{x}\in\mathbb{R}^{d}}F(\boldsymbol{x})=\frac{1}{n}\sum_{j=1}^{n}\left(\left|\langle\boldsymbol{\alpha}_{j},\boldsymbol{x}\rangle\right|-b_{j}\right)^{2}.$ (1.2)
To tackle model (1.2), Zhang et al. in [16] proposed the Reshaped Wirtinger Flow (RWF) through a simple gradient descent, while Wang et al. in [8] introduced Truncated Amplitude Flow (TAF) by truncated gradient descent; more recently Wang et al. in [17] designed Reweighted Amplitude Flow (RAF) via reweighted gradient descent. Interestingly, Zhang et al. in [16] proved that RWF enjoys the same sample complexity as TWF even without truncation in gradient steps.
Although numerical results indicate that algorithms based on the amplitude flow model (1.2) tend to outperform WF and TWF algorithms based on model (1.1) in both real and complex fields, the complex algorithms still lack rigorous theoretical guarantee. Several studies e.g., [33] have pointed out the theoretical analysis of RWF relies on the fact that the value of $\text{sign}\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle$ equals $-1$ or $1$ when $j=1,\cdots,n$, which can only be satisfied in the real number field. However, this condition is also essential for TWF, TAF and RAF, making straightforward generalization of theoretical results to the complex domain challenging. Some effective efforts have been made to address this problem, such as introducing new perturbed amplitude flow model [33].
Naturally, it is of interest to investigate the hypercomplex phase retrieval problem, where the signal $\boldsymbol{x}$ and the entries of the measurement matrix $\boldsymbol{A}$ in Prob. 1.1 take values from a specific hypercomplex algebra depending on the concrete applications. Indeed, recent interest has surged in hypercomplex phase retrieval (see, e.g., [4, 5, 9, 23]), although it is still in its infancy. Specifically, the quaternionic phase retrieval was first considered in [23] for real-valued sensing matrix, without giving a recovery algorithm. Moreover, the model in which $\boldsymbol{A}$ with real entries fails to harness the quaternion multiplication and leads to more trivial ambiguities that may limit applications of QPR. In [9], Chen and Ng considered quaternion-valued sensing matrix $\boldsymbol{A}\in\mathbb{H}^{n\times d}$, and introduced the quaternionic Wirtinger flow (QWF) algorithm, based on the generalized $\mathbb{HR}$ calculus and the quaternion version of the intensity-based model, i.e., $\mathbb{F}=\mathbb{H}$ in model (1.1). Additionally, counterparts of TWF and TAF in QPR were proposed and demonstrating numerical improvements over QWF. These algorithms were further specialized to pure quaternion signal, effectively applied to color image recovery. Notably, the quaternionic approach achieves reconstruction with significantly fewer phaseless measurements compared to traditional real-valued phase retrieval based on monochromatic model or concatenation model. However, akin to the complex $\mathbb{C}$ cases, QTWF and QTAF currently still lack theoretical guarantees. The octonionic phase retrieval is more challenging due to the non-associative and non-commutative structure of octonions, see [4].
1.2 Our Contribution
@@
In this work, we consider the quaternion phase retrieval problem and systemically investigate the quaternion non-convex phase retrieval algorithms based on the quaternionic version of the amplitude-based model (1.2). Note that one key advantage of quaternionic methods is their ability to recover signals with substantially fewer measurements compared to real-valued methods based on monochromatic or concatenation models. Our algorithms are not exception. Our contributions are two fold.
@@
(i) We introduce the Reweighted Amplitude Flow algorithm for quaternion-valued signals (QRAF), which extends the Real/Complex Amplitude Flow (RAF) algorithm originally proposed in [8]. Due to non-commutative nature of quaternion multiplication, the order of operations is crucial in our algorithm. Numerical results, presented in Section 7 show that QRAF algorithm consistently outperforms existing methods, such as QWF, QTWF and QTAF originally proposed in [9]. Notably, the Quaternionic Reshaped Wirtinger Flow (QRWF) and Quaternionic Truncated Amplitude Flow (QTAF) can be seen as special cases of QRAF. Furthermore, three variants of QRAF based on the gradient decent are introduced, which significantly enhance its performance. A detailed discussion for the convergence of QRAF is given, concluded with an open question which will be solved without significant efforts once it is solved in the complex setting.
@@
(ii) We also introduce the Quaternionic Perturbed Amplitude Flow (QPAF). The QPAF algprithm needn’t any truncation or re-weighted procedure, yet it achieves comparable numerical performance. Importantly, unlike the algorithms based on amplitude-based model (1.2), the theoretical analysis of QRAF is straightforward to extend to the present quaternion setting. Note that the QPAF has comparable recovery performance and computational efficiency with the QRAF.
@@
Paper Organization and Notations  The rest of this paper is organized as follows. Section 2 introduces the necessary preliminaries. In Section 3, we present the QRAF algorithm, and provide a discussion for its convergence. Section 4 explores three variants that further refine the QRAF algorithm. Section 5 focuses on the Quaternionic Perturbed Amplitude-based model and the non-convex QPAF algorithm. In Section 6, we review a useful technique—phase factor estimation—for color image processing. Experimental results on both synthetic and real-world color image data are discussed in Section 7. Concluding remarks are given in Section 8.
Throughout the paper, boldface lowercase letters such as $\boldsymbol{\alpha}_{i},\boldsymbol{x},\boldsymbol{z}$ denote vectors, and boldface capital letters such as $\boldsymbol{A},\boldsymbol{Y}$ denote matrices. For a quaternionic matrix and a vector in $\mathbb{H}^{d}$, while $\boldsymbol{A}^{*}$ and $\boldsymbol{z}^{*}$ denote conjugate transposes of $\boldsymbol{A}$ and $\boldsymbol{z}$, respectively. For a matrix and a vector, $\boldsymbol{A}^{T}$ and $\boldsymbol{z}^{T}$ denote transposes of $\boldsymbol{A}$ and $\boldsymbol{z}$, respectively.
@@ 2.1 Quaternionic Matrices
In mathematics, Hamilton’s quaternion algebra, denoted by $\mathbb{H}$, extends the familiar real and complex number fields. However, quternions do not form a field because their multiplication is generally non-commutative. A quaternion $q\in\mathbb{H}$ is typically represented as
$q=q_{a}+q_{b}i+q_{c}j+q_{d}k,$
where $q_{a},q_{b},q_{c},q_{d}\in\mathbb{R}$ and $i,j,k$ are generalized imaginary units satisfy the relations
$i^{2}=j^{2}=k^{2}=-1$
and
$ij=-ji=k,jk=-kj=i,ki=-ik=j.$
In this expression, the term $q_{b}i+q_{c}j+q_{d}k$ is called the vector part of $q$, while $q_{a}$ is referred to as the real or scalar part. The conjugate of $q$ is defined as $\bar{q}=q_{a}-q_{b}i-q_{c}j-q_{d}k$. For any two quaternions $p$ and $q$, then $\overline{pq}=\bar{q}\bar{p}$, which is in general not equal to $\bar{p}\bar{q}$ again due to the non-commutativity of quaternion multiplication. The Euclidean norm of $q$ is given by
$|q|=\sqrt{q\bar{q}}=\sqrt{q_{a}^{2}+q_{b}^{2}+q_{c}^{2}+q_{d}^{2}}.$
@@
The eigenvalue and eigenvector theory of a quaternion matrix is more complicated than those for real or complex matrices due to the non-commutativity. As a result, one can consider the left and right eigenvalue equations separately. In this work, we focus on the right eigenvalue and eigenvector which has its physical significance. Given ${\boldsymbol{A}}\in\mathbb{H}^{d\times d}$, if ${\boldsymbol{A}}{\bf x}={\boldsymbol{x}}\lambda$ for some nonzero ${\boldsymbol{x}}\in\mathbb{H}^{d}$, we refer $\lambda$, ${\boldsymbol{x}}$ as the right eigenvalue and eigenvector of ${\boldsymbol{A}}$. Note that ${\boldsymbol{A}\boldsymbol{x}}={\boldsymbol{x}}\lambda$ is equal to ${\boldsymbol{A}}({\boldsymbol{x}}v^{*})=(\boldsymbol{x})v^{*}(v\lambda v^{*})$ for any $v$ with modulus 1. Therefore, a matrix ${\boldsymbol{A}}$ with eigenvalue $\lambda$ has a set of eigenvalues $\{v\lambda v^{*}\}$, from which we can select a unique ‘standard eigenvalue’ in the form of $a+bi$ with $a\in\mathbb{R}$ and $b\geq 0)$. Any ${\boldsymbol{A}}\in\mathbb{H}^{d\times d}$ has exactly $d$ standard eigenvalues, and in particular, all standard eigenvalues of Hermitian ${\boldsymbol{A}}$ are real.
Similar to the eigenvalue decomposition for complex Hermitian matrices, a quaternion Hermitian matrix ${\boldsymbol{A}}$ can be decomposed as ${\boldsymbol{A}}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{U}^{*}$, where ${\boldsymbol{U}}$ is a unitary matrix and ${\boldsymbol{\Sigma}}$ is diagonal, with the standard eigenvalues of ${\boldsymbol{A}}$ arranged in the diagonal of ${\boldsymbol{\Sigma}}$. The detailed proof of the results mentioned above of quaternionic matrices can be found in [24].
@@ 2.2 Dirac Operator and Generalized $\mathbb{HR}$ Calculus
In this subsection, we review the fundamentals of quaternion matrix derivatives, specifically the generalized $\mathbb{HR}$ calculus introduced in [25]. This framework is essential for developing gradient-descent-like iterations in quaternionic phase retrieval algorithms. It provides a comprehensive set of rules that enable the computation of derivatives for functions directly within the quaternion domain, serving as an analogue to the Wirtinger calculus. It eliminates the need to convert the quaternionic optimization problem in (1.1) into an equivalent real-domain formulation.
For any quaternions $q$, consider the transformation
$q^{\mu}:=\mu q\mu^{-1}$
where $\mu$ is any non-zero quaternion, which represents a $3$-dimensional rotation of the vector part of $q$. The set $\{1,i^{\mu},j^{\mu},k^{\mu}\}$ forms a generalized orthogonal basis for $\mathbb{H}$. Similarly, for any quaternion vector $\boldsymbol{q}\in\mathbb{H}^{d}$, we define its transformation as $\boldsymbol{q}^{\mu}=\left(q_{1}^{\mu},q_{2}^{\mu},\cdots,q_{n}^{\mu}\right)$.
The generalized $\mathbb{HR}$ calculus can be derived either from left or right GHR derivatives. In this work, we will only use the left one. Note that in mathematical literature, the left derivative used here is often referred as the right derivative.
@@
The left generalized $\mathbb{HR}$ derivative of a function $f$ with respect to the transformed quaternion $q^{\mu}$ is defined as
$\frac{\partial_{\ell}f}{\partial q^{\mu}}=\frac{1}{4}\left(\frac{\partial f}{\partial q_{a}}-\frac{\partial f}{\partial q_{b}}i^{\mu}-\frac{\partial f}{\partial q_{c}}j^{\mu}-\frac{\partial f}{\partial q_{d}}k^{\mu}\right),$
@@
Remark 2.2.
The left and right GHR derivative operators are variants of the celebrated Dirac operators, which are widely studied in mathematical and physical literature, see e.g. [32].
@@
It can be seen that $\nabla_{\boldsymbol{q}}f(\boldsymbol{q})$ represents the direction of the steepest ascent of the scalar-valued function $f(\boldsymbol{q})$, indicating the direction of maximum rate of change.

### v2

# Quaternionic Reweighted Amplitude Flow for Phase Retrieval in Image Reconstruction

@@ Abstract
organization=Department of Electronics and Information Systems, Ghent University,addressline=Krijgslaan 281, Building S8, city=Gent, postcode=9000, state=Gent, country=Belgium
@@
organization=School of Mathematical Sciences, Tianjin Normal University,addressline=Binshui No.393, Xiqing District, city=Tianjin, postcode=300387, state=Tianjin, country=China
@@ 1 Introduction
Hypercomplex algebras, such as quaternions, octonions, and Clifford algebras, have demonstrated considerable utility in the representation and processing of multidimensional signals. These algebraic systems extend the capabilities of conventional vector spaces by enabling not only scalar multiplication and addition, but also non-commutative multiplication, allowing for compact and structured representations of color, geometric, and directional information. In particular, quaternion algebra has become increasingly important in color image processing [1], signal fusion [2] and neural networks [3], as it naturally encodes RGB channels in a single algebraic object, thereby preserving inter-channel correlations [2].
In recent years, quaternion-based methods have been increasingly applied to ill-posed inverse problems in imaging. Notably, a range of quaternion models have been introduced for color image denoising and inpainting. Quaternion Matrix Completion (QMC) [4] exploits the low-rank structure of quaternion-encoded images for recovering missing pixels; Nonlocal Self-Similarity QMC (NSS-QMC) [5] extends this by leveraging patch-based nonlocal priors to recover fine texture details; Saturation-Value Total Variation (SVTV) [6] applies a perceptual total variation regularization in the HSV color space, preserving color edges more faithfully; and Cross-Space Total Variation with Quaternion Blur Operator (CSTV-QBO) [7] introduces a novel model that simultaneously enforces TV constraints in multiple color spaces and models inter-channel blur using quaternion convolution. These methods demonstrate that quaternionic modeling not only enhances color consistency but also enables structure-preserving priors that are difficult to formulate in real-valued domains.
Inspired by this progress, researchers have recently begun extending classical phase retrieval (PR) to hypercomplex settings, see e.g., [8]. In the conventional PR problem, the objective is to recover a signal $\boldsymbol{x}$ from phaseless measurements of the form $b_{j}=|\langle\boldsymbol{\alpha}_{j},\boldsymbol{x}\rangle|$, see e.g. [9]. This inverse problem is fundamentally nonconvex and arises in various imaging applications such as optics, crystallography, and diffraction imaging [10]. Numerous algorithms have been proposed in the real and complex domains, including convex semi-definite relaxations [11], optimization-based method such as weighted nuclear norm minimization (WNNM) method [12], gradient-based methods such as Wirtinger Flow (WF) [13], and amplitude-based refinements such as Truncated Amplitude Flow (TAF) and Reweighted Amplitude Flow (RAF) [14].
However, extending these methods to the quaternionic domain introduces new challenges that cannot be trivially addressed. Unlike the complex field, the quaternion algebra is non-commutative, which complicates the gradient-based analysis. Moreover, many foundational tools—such as Wirtinger calculus and spectral theorem — require significant adaptation. Significant initial progress on quaternion phase retrieval (QPR) has been made by Chen and Ng in [15]. They proposed a Quaternion Wirtinger Flow (QWF) algorithm using a generalized $\mathbb{HR}$ calculus developed in [16], while the subsequent studies proposed Quaternion Truncated Wirtinger Flow (QTWF) and Quaternion Truncated Amplitude Flow (QTAF), which showed empirical improvements. More complicated Octonic phase retrieval was considered in [17]. Nevertheless, these methods primarily focused on intensity-based models.
1.1 Our Contributions
@@
In this work, we consider the quaternion phase retrieval problem and systemically investigate the quaternion non-convex phase retrieval algorithms based on the quaternionic amplitude-based model. Note that one key advantage of quaternionic methods is their ability to recover signals with substantially fewer measurements compared to real-valued methods based on monochromatic or concatenation models. Our algorithms are not exception. One of the primary difficulties in quaternionic phase retrieval stems from the non-commutative nature of quaternion multiplication, which complicates both algorithm design and theoretical analysis. Our proposed algorithms are designed to directly confront these issues. In particular, we leverage the quaternionic algebraic structure—rather than analyze it via real or complex reformulations—to develop efficient and theoretically motivated methods, thereby preserving inter-channel correlations. Our contributions are two fold:
@@
(i) We introduce the Reweighted Amplitude Flow algorithm for quaternion-valued signals (QRAF), which extends the Real/Complex Amplitude Flow (RAF) algorithm originally proposed in [18]. The non-commutativity of quaternion multiplication plays a central role in the algorithm’s design, necessitating careful ordering of operations to preserve correctness and convergence. Numerical experiments presented in Section 7 show that QRAF algorithm consistently outperforms existing methods, such as QWF, QTWF and QTAF [15]. Notably, both the Quaternionic Reshaped Wirtinger Flow (QRWF) and Quaternionic Truncated Amplitude Flow (QTAF) can be seen as special cases of QRAF. Furthermore, three variants of QRAF based on the gradient decent are introduced, which significantly enhance its performance. A detailed discussion for the convergence of QRAF is given, concluding with an open question which will be solved the corresponding complex-valued case is fully understood.
@@
(ii) We also propose the Quaternionic Perturbed Amplitude Flow (QPAF). The QPAF algorithm needn’t any truncation or re-weighted procedure, yet it achieves comparable numerical performance. Importantly, unlike the algorithms based on amplitude-based model, the theoretical analysis of PAF is straightforward to extend to the present quaternionic setting. making it a promising alternative in practice.
Since any quaternionic matrix can be represented using real or complex matrices, a natural and frequently asked question is that if our algorithms are merely rebranded versions of existing real/complex methods. In the context of quaternionic phase retrieval, this is not the case. Indeed, it is not possible to recover a signal $\boldsymbol{x}$ from phaseless quaternionic measurements of the form $b_{j}=|\langle\boldsymbol{\alpha}_{j},\boldsymbol{x}\rangle|$ using real or complex analogues alone. Our algorithms are intrinsically governed by the quaternionic algebraic structure. Even when implemented via real or complex matrix representations, the underlying quaternionic behavior remains critical to their performance and theoretical soundness.
@@
Paper Organization and Notations  The rest of this paper is organized as follows. Section 2 introduces the necessary preliminaries. In Section 3, we present the QRAF algorithm, and provide a discussion for its convergence. Section 4 explores three variants that further refine the QRAF algorithm. Section 5 focuses on the Quaternionic Perturbed Amplitude-based model and the non-convex QPAF algorithm. In Section 6, we review a useful technique—phase factor estimation—for color image processing. Experimental results on both synthetic and real-world color image data are discussed in Section 7. Concluding remarks are given in Section 8 and future works are discussed in 9. Throughout the paper, boldface lowercase letters such as $\boldsymbol{\alpha}_{i},\boldsymbol{x},\boldsymbol{z}$ denote vectors, and boldface capital letters such as $\boldsymbol{A},\boldsymbol{Y}$ denote matrices. For a quaternionic matrix and a vector in $\mathbb{H}^{d}$, while $\boldsymbol{A}^{*}$ and $\boldsymbol{z}^{*}$ denote conjugate transposes of $\boldsymbol{A}$ and $\boldsymbol{z}$, respectively. For a matrix and a vector, $\boldsymbol{A}^{T}$ and $\boldsymbol{z}^{T}$ denote transposes of $\boldsymbol{A}$ and $\boldsymbol{z}$, respectively.
@@ 2.1 Quaternionic Matrices
In mathematics, Hamilton’s quaternion algebra, denoted by $\mathbb{H}$, extends the familiar real and complex number fields. A quaternion $q\in\mathbb{H}$ is typically represented as $q=q_{a}+q_{b}i+q_{c}j+q_{d}k,$ where $q_{a},q_{b},q_{c},q_{d}\in\mathbb{R}$ and $i,j,k$ are generalized imaginary units satisfy the relations $i^{2}=j^{2}=k^{2}=-1$ and $ij=-ji=k,jk=-kj=i,ki=-ik=j$. In this expression, the term $q_{b}i+q_{c}j+q_{d}k$ is called the vector part of $q$, while $q_{a}$ is referred to as the real or scalar part. The conjugate of $q$ is defined as $\bar{q}=q_{a}-q_{b}i-q_{c}j-q_{d}k$. For any two quaternions $p$ and $q$, then $\overline{pq}=\bar{q}\bar{p}$, which is in general not equal to $\bar{p}\bar{q}$ again due to the non-commutativity of quaternion multiplication. The Euclidean norm of $q$ is given by $|q|=\sqrt{q\bar{q}}=\sqrt{q_{a}^{2}+q_{b}^{2}+q_{c}^{2}+q_{d}^{2}}.$
@@
The eigenvalue and eigenvector theory of a quaternion matrix is more complicated than those for real or complex matrices due to the non-commutativity. As a result, one can consider the left and right eigenvalue equations separately. In this work, we focus on the right eigenvalue and eigenvector which has its physical significance. Given ${\boldsymbol{A}}\in\mathbb{H}^{d\times d}$, if ${\boldsymbol{A}}{\bf x}={\boldsymbol{x}}\lambda$ for some nonzero ${\boldsymbol{x}}\in\mathbb{H}^{d}$, we refer $\lambda$, ${\boldsymbol{x}}$ as the right eigenvalue and eigenvector of ${\boldsymbol{A}}$. Note that ${\boldsymbol{A}\boldsymbol{x}}={\boldsymbol{x}}\lambda$ is equal to ${\boldsymbol{A}}({\boldsymbol{x}}v^{*})=(\boldsymbol{x})v^{*}(v\lambda v^{*})$ for any $v$ with modulus 1. Therefore, a matrix ${\boldsymbol{A}}$ with eigenvalue $\lambda$ has a set of eigenvalues $\{v\lambda v^{*}\}$, from which we can select a unique ‘standard eigenvalue’ in the form of $a+bi$ with $a\in\mathbb{R}$ and $b\geq 0)$. Any ${\boldsymbol{A}}\in\mathbb{H}^{d\times d}$ has exactly $d$ standard eigenvalues, and in particular, all standard eigenvalues of Hermitian ${\boldsymbol{A}}$ are real. Similar to complex Hermitian matrices, a quaternion Hermitian matrix ${\boldsymbol{A}}$ can be decomposed as ${\boldsymbol{A}}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{U}^{*}$, where ${\boldsymbol{U}}$ is a unitary matrix and ${\boldsymbol{\Sigma}}$ is diagonal, with the standard eigenvalues of ${\boldsymbol{A}}$ arranged in the diagonal of ${\boldsymbol{\Sigma}}$. The detailed proof of the results mentioned above of quaternionic matrices can be found in [19].
@@ 2.2 Dirac Operator and Generalized $\mathbb{HR}$ Calculus
In this subsection, we review the fundamentals of quaternion matrix derivatives. This framework is essential for developing gradient-descent-like iterations. It provides a comprehensive set of rules that enable the computation of derivatives for functions directly within the quaternion domain, serving as an analogue to the Wirtinger calculus. For any quaternions $q$, consider the transformation $q^{\mu}:=\mu q\mu^{-1}$ where $\mu$ is any non-zero quaternion, which represents a $3$-dimensional rotation of the vector part of $q$. The set $\{1,i^{\mu},j^{\mu},k^{\mu}\}$ forms a generalized orthogonal basis for $\mathbb{H}$. Similarly, for any quaternion vector $\boldsymbol{q}\in\mathbb{H}^{n}$, we define its transformation as $\boldsymbol{q}^{\mu}=\left(q_{1}^{\mu},q_{2}^{\mu},\cdots,q_{n}^{\mu}\right)$. The generalized $\mathbb{HR}$ calculus can be derived either from left or right GHR derivatives. In this work, we will only use the left one. Note that in mathematical literature (see e.g., [20]), the derivative used here is often referred as the right derivative.
@@
[16] The generalized $\mathbb{HR}$ derivative of a function $f$ with respect to the transformed quaternion $q^{\mu}$ is defined as
$\frac{\partial f}{\partial q^{\mu}}=\frac{1}{4}\left(\frac{\partial f}{\partial q_{a}}-\frac{\partial f}{\partial q_{b}}i^{\mu}-\frac{\partial f}{\partial q_{c}}j^{\mu}-\frac{\partial f}{\partial q_{d}}k^{\mu}\right),$
@@
It can be seen that $\nabla_{\boldsymbol{\bar{q}}}f(\boldsymbol{q})$ represents the direction of the steepest ascent of the scalar-valued function $f(\boldsymbol{q})$, indicating the direction of maximum rate of change, see e.g., [21].